suppose demand and total cost functions for a monopoly firm are given as follow,
demand function- p=800-3Q
Cost function- Tc=60+40Q+Q2
requered;-
1. total revenue function
2.average revenue function
3. average cost function
4.profit maximizing output
5.profit maximizing price
1
Expert's answer
2020-06-29T15:07:10-0400

1) Total revenue function.

"=Price(quantity)"

"=(800-3Q)Q"

"=800Q-3Q"2

2) Average revenue function

Average revenue function"="

"=\frac{\text{Total revenue}}{{Quantity}}"


"=\frac{800Q-3Q^{2}}{Q}"

"\text{Average revenue function}=800-3Q"

3) Average cost function

"\text{Average cost function}="

"\frac{\text{Total cost function}}{\text{Quantity}}"

"\text{Average cost function}="

"\frac{60+40Q+Q^2}{Q}"

"\text{Average cost function}" "="

"\frac{60}{Q}+40+Q"


4) Profit maximizing output

"\text{Profit maximizing output}="

"MR=MC"

"MR=\frac{\text{∂TR}}{\text{∂Q}}"

"=800Q-3Q^2"

"=800-6Q"

"MC=\frac{∂TC}{∂Q}"

"=60+40Q+Q^2"

"=40+2Q"

"=MR=MC"

"=800-6Q=40+2Q"

"=800-40=2Q+6Q"

"760=8Q"

"Q=95"

Profit is maximized at "Q=95"

5) Profit maximizing price

"P=800-3Q"

"P=800-3(95)"

"P=800-285"

"P=515"

"\text{Profit maximizing price is}" "515"